rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#12849 — FS#17728 — Lyon optical loop

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
Following a performance drop of signal on the optical loop of Lyon, we will
work on the lal69 site to replace an attenuator.
Date:  Tuesday, 26 April 2016, 19:54PM
Reason for closing:  Done
Comment by OVH - Tuesday, 26 April 2016, 19:50PM

A technician is on site.


Comment by OVH - Tuesday, 26 April 2016, 19:51PM

Intervention complete. We have replaced attenuator 2dB by 10dB.

We will now work on mpl69.


Comment by OVH - Tuesday, 26 April 2016, 19:53PM

We have removed -5dB attenuator on the mpl69 site.

No more CRC are detected in the loop.